OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)?

OpenStudy (anonymous):

oh thats easy...just plug it in to y=mx+b i think

OpenStudy (johnweldon1993):

Well first...we need to find slope you have 2 points (4, -1) (-2, 3) (x1,y1)(x2,y2) plug those (using the labeling I provided below the points) into this equation \[m = \frac{ y2 - y1 }{ x2 - x1 }\] This will be your slope of the line

OpenStudy (anonymous):

-2x + 3y - 5 = 0 is this correct

OpenStudy (johnweldon1993):

not quite...you can always check your answer by plugging in a point that you were given and checking if it works...in that case -2(4) + 3(-1) - 5 = 0 -8 -3 -5 =0 -16 = 0 Doesn't work... Try it the way I have explained above...by finding the slope first

OpenStudy (anonymous):

m=4/-2 ?

OpenStudy (johnweldon1993):

close 3 - (-1) 4 ------- = ------ -2 - 4 -6 When simplified becomes -2/3 right?

OpenStudy (anonymous):

that's what i had the first time but it didn't seem right, but i guess it is

OpenStudy (johnweldon1993):

never second guess yourself :) but okay....That is your 'm' Now, pick which of the 2 point you were given...that you would like to use...

OpenStudy (anonymous):

4,-1

OpenStudy (johnweldon1993):

okay Using this equation y = mx + b 'y' will be the y-value in the point you chose 'x' will be the x-value in the point you chose 'm' will be the slope we just found so after plugging everything in -1 = (-2/3)(4) + b Solve for b

OpenStudy (anonymous):

is B, the next point?

OpenStudy (johnweldon1993):

'b' is what we're solving for...and this will be the y-intercept of your lines equation...Once you solve for 'b' here...you will have your full equation as you will see...

OpenStudy (anonymous):

to solve you do multiply -2/3 and 4

OpenStudy (johnweldon1993):

Yes you can start there....after you do that you will have -1 = (-8/3) + b What would we do next?

OpenStudy (anonymous):

would it be -2 or 2

OpenStudy (johnweldon1993):

not quite next you would add (8/3) to both sides to cancel it out on the right side -1 + (8/3) = b finish simplifying that left side what is -1 + (8/3)?

OpenStudy (anonymous):

7/3

OpenStudy (johnweldon1993):

not quite....you need to find a least common denominator when adding or subtracting fractions the lcd here...would be 3...so you need to multiply -1 by (3/3) this will give you -3 8 --- + --- 3 3 now you can combine the fractions and you'll get -3 + 8 ------ 3 what does this simplify to?

OpenStudy (anonymous):

5/3

OpenStudy (johnweldon1993):

perfect....that is your y-intercept Now....you have everything you need for your line equation you have a slope (-2/3) and you have a y-intercept (5/3) putting that all together you get \[y = -\frac{ 2 }{ 3 }x + \frac{ 5 }{ 3 }\] That would be the equation of your line

OpenStudy (anonymous):

THanks!

OpenStudy (johnweldon1993):

No problem! I hope it was understandable?